# On the metric of the jet bundle and similarity on weighted Dirichlet spaces

### v2

@@ 4. Similarity on weighted Dirichlet spaces
Finally, by combining $\mathcal{E}_{T}\sim_{u}\mathcal{E}_{D_{\alpha}^{*}}\otimes\mathcal{E}_{D_{T}}$, we conclude that $T\sim_{s}D^{*}_{\alpha}$. ∎
In Theorem 4.3, the ratio of the metrics of holomorphic line bundles $\mathcal{E}_{T}$ and $\mathcal{E}_{D^{*}_{\alpha}}$ is the metric of the holomorphic line bundle $\mathcal{E}_{D_{T}}$ induced by the defect operator $D_{T}$, that is, $h_{\mathcal{E}_{D_{T}}}=\frac{h_{\mathcal{E}_{T}}}{h_{D^{*}_{\alpha}}}$. Letting $\mathcal{H}_{K}$ is a Hilbert space with diagonal reproducing kernel ${K}(z,w)=\sum\limits_{m,n=0}^{\infty}a_{m,n}z^{m}\overline{w}^{n}$, which means that for all $n\neq m$, $a_{m,n}=0$. From this, the similarity classification of holomorphic complex bundles and the similarity of backward shift operators on non-diagonal reproducing kernel Hilbert space can be obtained.
@@
Let $\mathcal{E}_{1}$ and $\mathcal{E}_{2}$ be Hermitian holomorphic line bundle over $\mathbb{D}$ and $\mathcal{E}_{1}\sim_{u}\mathcal{E}_{D_{\alpha}^{*}}\otimes\mathcal{E}_{2}$. If there exist constants $c_{2}>c_{1}>0$ and a metric $h_{\mathcal{E}_{2}}$ of $\mathcal{E}_{2}$ satisfy
$c_{1}\leq h_{\mathcal{E}_{2}}(w)<\text{trace}\;\mathcal{J}_{1}(h_{\mathcal{E}_{2}})(w)\leq c_{2}$
@@
Let $T\in\mathcal{B}_{1}(\mathbb{D})$ and $\mathcal{E}_{T}\sim_{u}\mathcal{E}_{D_{\alpha}^{*}}\otimes\mathcal{E}$ for some Hermitian holomorphic line bundle $\mathcal{E}$ over $\mathbb{D}$. If the metric $h_{\mathcal{E}}$ of $\mathcal{E}$ satisfies $h_{\mathcal{E}}(w)=1+\sum\limits_{i=0}^{n}|\phi_{i}(w)|^{2}$ for polynomials $\phi_{1},\ldots,\phi_{n}$, and positive integer $n$, then $\mathcal{E}_{T}\sim_{s}\mathcal{E}_{D_{\alpha}^{*}}$, that is, $T\sim_{s}D_{\alpha}^{*}$.
Given the discussion in this section, we use holomorphic line bundles to provide a necessary and sufficient for certain operators to be similar to backward shift operators on weighted Dirichlet spaces.
@@
If the reproducing kernel of a Hilbert space $\mathcal{H}_{K}$ is of the form ${K}(z,w)=\sum\limits_{m,n=0}^{\infty}a_{m,n}z^{m}\overline{w}^{n},$ where, for some $n\neq m$, $a_{m,n}\neq 0$, then Theorem 2.7 cannot be used to determine similarity. The following example is a good illustration of how one can use Theorem 4.3 to conclude similarity:
@@
Let $M_{z}^{*}$ be the backward shift operator on Hilbert space $\mathcal{H}_{K}$ with reproducing kernel
@@
Since the reproducing kernel of Dirichlet space $\mathcal{D}$ is $K_{1}(z,w)=\sum\limits_{i=0}^{\infty}\frac{z^{i}\overline{w}^{i}}{i+1}$, $z,w\in\mathbb{D}$, we obtain that $K(z,w)=K_{1}(z,w)\cdot\left[1+(1+z+z^{2})(1+\overline{w}+\overline{w}^{2})\right]$. Then there exists a Hermitian holomorphic line bundle $\mathcal{E}$ over $\mathbb{D}$ such that $\mathcal{E}_{M_{z}^{*}}=\mathcal{E}_{D^{*}}\otimes\mathcal{E}$, and a non-vanishing holomorphic cross-section $\gamma$ of $\mathcal{E}$ satisfies $\|\gamma(w)\|^{2}=1+|1+w+w^{2}|^{2}$. Therefore, the metric $h_{\mathcal{E}}(w)=\|\gamma(w)\|^{2}$ of $\mathcal{E}$ satisfies
$1\leq h_{\mathcal{E}}(w)<\text{trace}\mathcal{J}_{1}(h_{\mathcal{E}})(w)\leq 19,\quad w\in\mathbb{D}.$
Hence the conclusion can be derived from Theorem 4.3. ∎
In Theorem 4.3, the condition that $\mathcal{E}_{D_{T}}(w)\subset\mathbb{C}^{n}$ for some integer $n>0$ was necessary. A natural question is whether the result also exists when $n=\infty$. The following example answers the question in the affirmative:
@@
Let $M_{z}^{*}$ be the backward shift operator on Hilbert space $\mathcal{H}_{{K}}$ with reproducing kernel
@@
From the relationship between the weight sequence and the reproducing kernel of Hilbert space, we obtain that the weight sequence of $M_{z}^{*}$ is $\left\{\alpha_{n}=\sqrt{\frac{a_{n-1}}{a_{n}}}\right\}_{n\geq 1}$, where $a_{n}=\sum\limits_{i=0}^{n}\frac{1}{(i+1)(n-i+1)^{2}}.$ Since the weight sequence of $D^{*}$ is $\left\{\beta_{n}=\sqrt{\frac{n+1}{n}}\right\}_{n\geq 1}$, we have that $\prod\limits_{i=1}^{n}\beta_{i}\Big{/}\prod\limits_{i=1}^{n}\alpha_{i}=\sqrt{(n+1)a_{n}}\geq 1.$
Now, to prove from Theorem 2.7 that $M_{z}^{*}$ is similar to $D^{*}$, we only need to prove that $\lim\limits_{n\rightarrow\infty}(n+1)a_{n}=c$ for some constant $c.$
@@
Next is another major theorem in this section, as Theorem 4.3, we provide another criterion for the similarity of backward shift operators in the weighted Dirichlet space when $n=\infty$.
@@
Let $T\in\mathcal{B}_{1}(\mathbb{D})$ and $\sum\limits_{s=1}^{\infty}\frac{\|T^{s}\|^{2}}{(s+1)^{\alpha}}\leq 1$ for some constant $\alpha\in(0,1]$. If there is a non-vanishing holomorphic cross-section $\gamma$ of $\mathcal{E}_{D_{T}}$ such that for some $\delta>0$,
@@
Now let $F^{\#}(w):=F(\overline{w})$ and hence, $(F^{\#})^{*}\in\text{Mult}(\mathcal{D}_{\alpha,E},\mathcal{D}_{\alpha})$. If we denote by $M_{\gamma}^{C}$ the column operator from $\mathcal{D}_{\alpha}$ to $\bigoplus\limits_{1}^{\infty}\mathcal{D}_{\alpha}$ defined by
@@
is bounded and again from Theorem 2.3, we know that the row operator $M_{G}^{R}:\bigoplus\limits_{1}^{\infty}\mathcal{D}_{\alpha}\rightarrow\mathcal{D}_{\alpha}$ is bounded as well.
@@
The condition $\sum\limits_{s=1}^{\infty}\frac{\|T^{s}\|^{2}}{(s+1)^{\alpha}}\leq 1$ in Theorem 4.9 ensures that the Hermitian holomorphic line bundle $\mathcal{E}_{T}$ acquires a tensor product structure. The natural question arises, what type of operator corresponds to a Hermitian holomorphic complex bundle with a tensor product structure in the unitarily equivalent sense? We will give a precise answer, distinct from the condition $\sum\limits_{s=1}^{\infty}\frac{\|T^{s}\|^{2}}{s+1}\leq 1$, through the following example.
@@
Let $M_{z}^{*}$ be the backward shift operator on the Hilbert space $\mathcal{H}_{K}$ with reproducing kernel $K(z,w)=\sum\limits_{n=0}^{\infty}a_{n}z^{n}\overline{w}^{n}$, $z,w\in\mathbb{D}$, where $a_{n}=\sum\limits_{\mbox{\tiny$\begin{array}[]{c}i+j^{5}=n\\ i,j\in\mathbb{N}\end{array}$}}\frac{1}{(i+1)(j+1)^{8}}$ for $n\in\mathbb{N}$. We have that $\left\{\sqrt{a_{n}}z^{n}\right\}_{n\geq 0}$ is an orthonormal basis of $\mathcal{H}_{K}$ and $\left\{\sqrt{\frac{a_{n-1}}{a_{n}}}\right\}_{n\geq 1}$ is the weight sequence of $M_{z}^{*}$. After calculations, we obtain that
@@
It follows that $\sum\limits_{s=1}^{\infty}\frac{M_{z}^{s}M_{z}^{*s}}{s+1}\leq I$, and then $a_{n}-\frac{1}{2}a_{n-1}-\cdots-\frac{1}{n+1}a_{0}\geq 0$ for any positive integer $n$. However, for $a_{2}=\frac{515}{1536}$, $a_{1}=\frac{129}{256}$ and $a_{0}=1$, we have that $a_{2}-\frac{1}{2}a_{1}-\frac{1}{3}a_{0}=\frac{515}{1536}-\frac{1}{2}\times\frac{129}{256}-\frac{1}{3}=-\frac{1}{4}<0$. This means that $M_{z}^{*}$ does not satisfy $\sum\limits_{s=1}^{\infty}\frac{\|M_{z}^{*s}\|^{2}}{s+1}\leq 1$.
Setting $\mathcal{E}$ is a Hermitian holomorphic line bundle over $\mathbb{D}$ with a non-vanishing holomorphic cross-section $\gamma$ satisfying $\|\gamma(w)\|^{2}=\sum\limits_{j=0}^{\infty}\frac{|w|^{2j^{5}}}{(j+1)^{8}}$, then from (4.6), $\mathcal{E}_{M_{z}^{*}}\sim_{u}\mathcal{E}_{D^{*}}\otimes\mathcal{E}$ for the backward shift operator $D^{*}$ on the Dirichlet space $\mathcal{D}$. Let $E=\bigvee\limits_{w\in\mathbb{D}}\gamma(w)$ and $\{e_{k}\}_{k=0}^{\infty}$ be an orthonormal basis of $E$, then $\gamma(w)=\sum\limits_{k=0}^{\infty}\gamma_{k}(w)e_{k}$ for $\gamma_{k}(w)=\frac{w^{k^{5}}}{(k+1)^{4}}$. For any holomorphic function $f\in\mathcal{D}$ with Taylor expansion $f(z)=\sum\limits_{n=0}^{\infty}\hat{f}(n)z^{n}$, there is $\|f\|_{\mathcal{D}}^{2}=\sum\limits_{n=0}^{\infty}(n+1)|\hat{f}(n)|^{2}<\infty$. Therefore, for $\gamma_{k}f,k\geq 0$, we have that $\gamma_{k}(z)f(z)=\sum\limits_{n=0}^{\infty}\frac{\hat{f}(n)}{(k+1)^{4}}z^{n+k^{5}}$, and then
@@
$\leq$ $\left[\frac{1}{(k+1)^{8}}+\frac{1}{(k+1)^{5}}\right]\|f\|_{\mathcal{D}}^{2}.$
It follows that $\|\gamma_{k}\|^{2}_{\text{Mult}(\mathcal{D})}\leq\frac{1}{(k+1)^{8}}+\frac{1}{(k+1)^{5}}\leq\frac{2}{(k+1)^{2}}$, Hence,
@@
Using the proof of Theorem 4.9, we obtain that $M_{z}^{*}\sim_{s}D^{*}$.
From Example 4.10, we demonstrate that the weighted backward shift operator $M_{z}^{*}$ does not satisfy the condition $\sum\limits_{s=1}^{\infty}\frac{\|M_{z}^{*s}\|^{2}}{s+1}\leq 1$. Furthermore, using Shields’ result, Theorem 2.7, it is not straightforward to verify that $M_{z}^{*}$ is similar to $D^{*}$. Now, let’s consider an example of a Hilbert space with a non-diagonal reproducing kernel. The proof, being similar to that of Example 4.10, is omitted here.
@@
Let $M_{z}^{*}$ be the backward shift operator on Hilbert space $\mathcal{H}_{K}$ with reproducing kernel
@@
We close this section with a proposition that shows that if a Hermitian holomorphic complex bundle with a tensor product satisfies the given conditions, there exists a corresponding bounded linear operator in $\mathcal{B}_{1}(\mathbb{D})$. This establishes the operator realization of the tensor product of holomorphic complex bundles.
@@
Let $\mathcal{E}$ be a Hermitian holomorphic line bundle over $\mathbb{D}$, and let $\gamma$ be a non-vanishing holomorphic cross-section of $\mathcal{E}$. Suppose that $E=\bigvee\limits_{w\in\mathbb{D}}\gamma(w),$ and $\{e_{i}\}_{i=0}^{\infty}$ is an orthonormal basis of $E$. If
@@
for $\gamma(w)=\sum\limits_{i=0}^{\infty}\gamma_{i}(w)e_{i}$ and some constant $\delta>0$, then there exists operator $T\in\mathcal{B}_{1}(\mathbb{D})$ such that $\mathcal{E}_{T}=\mathcal{E}_{D_{\alpha}^{*}}\otimes\mathcal{E}$ and $T\sim_{s}D_{\alpha}^{*}.$
@@
Recall the Cowen-Douglas conjecture is that for operators $T$ and $S$ in $\mathcal{B}_{1}(\mathbb{D})$ having $\overline{\mathbb{D}}$ as a $K$-spectral set, $T$ is similar to $S$ if and only if $\lim\limits_{|w|\rightarrow 1}\mathcal{K}_{T}(w)/\mathcal{K}_{S}(w)=1$. In [10, 11], Clark and Misra constructed examples to show that the Cowen-Douglas conjecture is false and claimed that the ratio of the metrics can be a better indication of similarity than the ratio of the curvatures. Meanwhile, supporting example are provided in Theorem 1 of [11]. Theorem 3.4 and Proposition 3.6 in this note are positive results supporting the claim made by Clark and Misra. Then a natural question is, which Cowen-Douglas operator’s similarity invariant is the ratio of the metrics?
Every Cowen-Douglas operator is unitary equivalent to the adjoint of a multiplication operator on a reproducing kernel Hilbert space consisting of holomorphic functions. For Hardy space and weighted Bergman spaces, Theorem 1 in [11], Theorem 3.8 in [55], Theorem 3.4 in [25], Theorem 3.4 and Proposition 3.6 suggest that, under certain conditions, the ratio of the metrics is a similarity invariant of the backward shift operator on these spaces. In particular, for the $n$-hypercontraction $T\in\mathcal{B}_{1}(\mathbb{D})$, from Theorem 0.1 in [38], Theorem 2.4 in [16], Theorem 2.3 in [24], and Theorem 3.4, we obtain that the operator $T$ is similar to $S_{n}^{*}$ on reproducing kernel Hilbert space $\mathcal{M}_{n}$ if and only if $a_{w}:=h_{\mathcal{E}_{T}}(w)/h_{\mathcal{E}_{S_{n}^{*}}}(w)$ is bonded and bounded from 00 if and only if the curvature $\mathcal{K}_{\mathcal{E}_{D_{T}}}$ of the holomorphic line bundle $\mathcal{E}_{D_{T}}$ induced by the defect operator $D_{T}$ satisfies $\mathcal{K}_{\mathcal{E}_{D_{T}}}=-\frac{\partial^{2}}{\partial w\partial\overline{w}}\varphi(w)$ for some bounded subharmonic function $\varphi$ defined on $\mathbb{D}$.
Clark and Misra constructed two contraction backward shift operators in 2.5 of [11], such that the ratio of the metrics of the holomorphic line bundles associated with operators is bounded and bounded from 00, but they are not similar. This indicates that the ratio of metrics is not a similarity invariant for all backward shift operators on the Hilbert space of diagonal reproducing kernels. However, the following question is unknown:

### v3

@@ 4. Similarity on weighted Dirichlet spaces
Finally, by considering $\mathcal{E}_{T}\sim_{u}\mathcal{E}_{D_{\alpha}^{*}}\otimes\mathcal{E}_{D_{T}}$, we conclude that $T\sim_{s}D^{*}_{\alpha}$. ∎
In Theorem 4.3, the ratio of the metrics of holomorphic line bundles $\mathcal{E}_{T}$ and $\mathcal{E}_{D^{*}_{\alpha}}$ is the metric of the holomorphic line bundle $\mathcal{E}_{D_{T}}$ induced by the defect operator $D_{T}$, that is, $h_{\mathcal{E}_{D_{T}}}=\frac{h_{\mathcal{E}_{T}}}{h_{D^{*}_{\alpha}}}$. Letting $\mathcal{H}_{K}$ be a Hilbert space with a diagonal reproducing kernel ${K}(z,w)=\sum\limits_{m,n=0}^{\infty}a_{m,n}z^{m}\overline{w}^{n}$, which means that for all $n\neq m$, $a_{m,n}=0$, we obtain a similarity classification of holomorphic complex bundles and the similarity of backward shift operators on non-diagonal reproducing kernel Hilbert spaces.
@@
Let $\mathcal{E}_{1}$ and $\mathcal{E}_{2}$ be Hermitian holomorphic line bundles over $\mathbb{D}$ where $\mathcal{E}_{1}\sim_{u}\mathcal{E}_{D_{\alpha}^{*}}\otimes\mathcal{E}_{2}$. If there exist constants $c_{2}>c_{1}>0$ and a metric $h_{\mathcal{E}_{2}}$ of $\mathcal{E}_{2}$ such that
$c_{1}\leq h_{\mathcal{E}_{2}}(w)<\text{trace}\;\mathcal{J}_{1}(h_{\mathcal{E}_{2}})(w)\leq c_{2},$
@@
Let $T\in\mathcal{B}_{1}(\mathbb{D})$ and let $\mathcal{E}_{T}\sim_{u}\mathcal{E}_{D_{\alpha}^{*}}\otimes\mathcal{E}$ for some Hermitian holomorphic line bundle $\mathcal{E}$ over $\mathbb{D}$. If the metric $h_{\mathcal{E}}$ of $\mathcal{E}$ satisfies $h_{\mathcal{E}}(w)=1+\sum\limits_{i=1}^{n}|\phi_{i}(w)|^{2}$ for some polynomials $\phi_{1},\ldots,\phi_{n}$, and a positive integer $n$, then $\mathcal{E}_{T}\sim_{s}\mathcal{E}_{D_{\alpha}^{*}}$, that is, $T\sim_{s}D_{\alpha}^{*}$.
A necessary and sufficient condition for certain operators to be similar to the backward shift operators on weighted Dirichlet spaces in terms of holomorphic line bundles is given next as follows:
@@
If the reproducing kernel of a Hilbert space $\mathcal{H}_{K}$ is of the form ${K}(z,w)=\sum\limits_{m,n=0}^{\infty}a_{m,n}z^{m}\overline{w}^{n},$ where, for some $n\neq m$, $a_{m,n}\neq 0$, then Theorem 2.7 cannot be used to determine similarity. The following example is a good illustration of how one can use Theorem 4.3 to conclude that two operators are similar:
@@
Let $M_{z}^{*}$ be the backward shift operator on a Hilbert space $\mathcal{H}_{K}$ with reproducing kernel
@@
Since the reproducing kernel of the Dirichlet space $\mathcal{D}$ is given as $K_{1}(z,w)=\sum\limits_{i=0}^{\infty}\frac{z^{i}\overline{w}^{i}}{i+1}$, $z,w\in\mathbb{D}$,
$K(z,w)=K_{1}(z,w)\cdot\left[1+(1+z+z^{2})(1+\overline{w}+\overline{w}^{2})\right].$
Then there exists a Hermitian holomorphic line bundle $\mathcal{E}$ over $\mathbb{D}$ such that $\mathcal{E}_{M_{z}^{*}}=\mathcal{E}_{D^{*}}\otimes\mathcal{E}$. Moreover, a non-vanishing holomorphic cross-section $\gamma$ of $\mathcal{E}$ is such that $\|\gamma(w)\|^{2}=1+|1+w+w^{2}|^{2}$. Therefore, the metric $h_{\mathcal{E}}(w)=\|\gamma(w)\|^{2}$ of $\mathcal{E}$ satisfies
$1\leq h_{\mathcal{E}}(w)<\text{trace}\mathcal{J}_{1}(h_{\mathcal{E}})(w)\leq 19,\quad w\in\mathbb{D},$
and the conclusion can be derived from Theorem 4.3. ∎
In Theorem 4.3, the condition that $\mathcal{E}_{D_{T}}(w)\subset\mathbb{C}^{n}$ for some integer $n>0$ was necessary. A natural question is whether the result also holds when $n=\infty$. The following example answers the question in the affirmative:
@@
Let $M_{z}^{*}$ be the backward shift operator on a Hilbert space $\mathcal{H}_{{K}}$ with reproducing kernel
@@
Using the relationship between a weight sequence and the reproducing kernel of a Hilbert space, we see that the weight sequence of $M_{z}^{*}$ is $\left\{\alpha_{n}=\sqrt{\frac{a_{n-1}}{a_{n}}}\right\}_{n\geq 1}$, where $a_{n}=\sum\limits_{i=0}^{n}\frac{1}{(i+1)(n-i+1)^{2}}.$ Since the weight sequence of $D^{*}$ is $\left\{\beta_{n}=\sqrt{\frac{n+1}{n}}\right\}_{n\geq 1}$, we have that $\prod\limits_{i=1}^{n}\beta_{i}\Big{/}\prod\limits_{i=1}^{n}\alpha_{i}=\sqrt{(n+1)a_{n}}\geq 1.$
@@
The following is another major theorem of this section. We provide a criterion for the similarity of backward shift operators on weighted Dirichlet spaces when $n=\infty$:
@@
Let $T\in\mathcal{B}_{1}(\mathbb{D})$ and suppose that $\sum\limits_{s=1}^{\infty}\frac{\|T^{s}\|^{2}}{(s+1)^{\alpha}}\leq 1$ for some constant $\alpha\in(0,1]$. If there is a non-vanishing holomorphic cross-section $\gamma$ of $\mathcal{E}_{D_{T}}$ such that for some $\delta>0$,
@@
Now let $F^{\#}(w):=F(\overline{w})$. Obviously, $(F^{\#})^{*}\in\text{Mult}(\mathcal{D}_{\alpha,E},\mathcal{D}_{\alpha})$. If we denote by $M_{\gamma}^{C}$ the column operator from $\mathcal{D}_{\alpha}$ to $\bigoplus\limits_{1}^{\infty}\mathcal{D}_{\alpha}$ defined by
@@
is bounded and from Theorem 2.3, we know that the row operator $M_{G}^{R}:\bigoplus\limits_{1}^{\infty}\mathcal{D}_{\alpha}\rightarrow\mathcal{D}_{\alpha}$ is bounded as well.
@@
Remark 4.10.
The condition (4.5) in Theorem 4.9 is a sufficient but not necessary condition, as illustrated in Example 4.8. In Example 4.8, $M_{z}^{*}$ is similar to $D^{*}$. From the holomorphic frame $\gamma$ of the holomorphic vector bundle $\mathcal{E}$, we obtain $\gamma(w)=\sum\limits_{i=1}^{\infty}\gamma_{i}(w)e_{i}$, where $\gamma_{i}(w)=\frac{1}{i}w^{i-1}$, and hence
$\sum\limits_{i=1}^{\infty}\|\gamma_{i}\|^{2}_{\text{Mult}(\mathcal{D})}\geq\sum\limits_{i=1}^{\infty}\|\gamma_{i}\|_{\mathcal{D}}^{2}=\sum\limits_{i=1}^{\infty}\frac{1}{i}=\infty.$
The condition $\sum\limits_{s=1}^{\infty}\frac{\|T^{s}\|^{2}}{(s+1)^{\alpha}}\leq 1$ in Theorem 4.9 ensures that the Hermitian holomorphic line bundle $\mathcal{E}_{T}$ acquires a tensor product structure. A natural question then arises: which types of operators correspond to Hermitian holomorphic complex bundles with a tensor product structure in the sense of unitary equivalence? We will give the precise answer, distinct from the condition $\sum\limits_{s=1}^{\infty}\frac{\|T^{s}\|^{2}}{s+1}\leq 1$, through the following example:
@@
Let $M_{z}^{*}$ be the backward shift operator on a Hilbert space $\mathcal{H}_{K}$ with reproducing kernel $K(z,w)=\sum\limits_{n=0}^{\infty}a_{n}z^{n}\overline{w}^{n}$, $z,w\in\mathbb{D}$, where $a_{n}=\sum\limits_{\mbox{\tiny$\begin{array}[]{c}i+j^{5}=n\\ i,j\in\mathbb{N}\end{array}$}}\frac{1}{(i+1)(j+1)^{8}}$ for $n\in\mathbb{N}$. In this case, $\left\{\sqrt{a_{n}}z^{n}\right\}_{n\geq 0}$ is an orthonormal basis for $\mathcal{H}_{K}$ and $\left\{\sqrt{\frac{a_{n-1}}{a_{n}}}\right\}_{n\geq 1}$ is the weight sequence of $M^{*}_{z}$. It can be computed that
@@
It follows that $\sum\limits_{s=1}^{\infty}\frac{M_{z}^{s}M_{z}^{*s}}{s+1}\leq I$, so that $a_{n}-\frac{1}{2}a_{n-1}-\cdots-\frac{1}{n+1}a_{0}\geq 0$ for every positive integer $n$. However, for $a_{2}=\frac{515}{1536}$, $a_{1}=\frac{129}{256}$ and $a_{0}=1$, $a_{2}-\frac{1}{2}a_{1}-\frac{1}{3}a_{0}=\frac{515}{1536}-\frac{1}{2}\times\frac{129}{256}-\frac{1}{3}=-\frac{1}{4}<0$. This means that $M_{z}^{*}$ does not satisfy $\sum\limits_{s=1}^{\infty}\frac{\|M_{z}^{*s}\|^{2}}{s+1}\leq 1$.
If we set $\mathcal{E}$ to be a Hermitian holomorphic line bundle over $\mathbb{D}$ with a non-vanishing holomorphic cross-section $\gamma$ satisfying $\|\gamma(w)\|^{2}=\sum\limits_{j=0}^{\infty}\frac{|w|^{2j^{5}}}{(j+1)^{8}}$, then from (4.6), $\mathcal{E}_{M_{z}^{*}}\sim_{u}\mathcal{E}_{D^{*}}\otimes\mathcal{E}$ for the backward shift operator $D^{*}$ on the Dirichlet space $\mathcal{D}$. Let $E=\bigvee\limits_{w\in\mathbb{D}}\gamma(w)$ and let $\{e_{k}\}_{k=0}^{\infty}$ be an orthonormal basis of $E$, so that $\gamma(w)=\sum\limits_{k=0}^{\infty}\gamma_{k}(w)e_{k},$ where $\gamma_{k}(w)=\frac{w^{k^{5}}}{(k+1)^{4}}$.
Since for $f\in\mathcal{D}$, $\|f\|_{\mathcal{D}}^{2}=\sum\limits_{n=0}^{\infty}(n+1)|\hat{f}(n)|^{2}<\infty$, it is easily seen that $\gamma_{k}(z)f(z)=\sum\limits_{n=0}^{\infty}\frac{\hat{f}(n)}{(k+1)^{4}}z^{n+k^{5}},k\geq 0$. It follows that
@@
$\leq$ $\left[\frac{1}{(k+1)^{8}}+\frac{1}{(k+1)^{3}}\right]\|f\|_{\mathcal{D}}^{2}.$
Therefore, $\|\gamma_{k}\|^{2}_{\text{Mult}(\mathcal{D})}\leq\frac{1}{(k+1)^{8}}+\frac{1}{(k+1)^{3}}\leq\frac{2}{(k+1)^{2}}$, and hence,
@@
Using the proof of Theorem 4.9, we conclude that $M_{z}^{*}\sim_{s}D^{*}$.
In Example 4.11, it was shown that the weighted backward shift operator $M_{z}^{*}$ does not satisfy the condition $\sum\limits_{s=1}^{\infty}\frac{\|M_{z}^{*s}\|^{2}}{s+1}\leq 1$. Furthermore, it is not straightforward to verify that $M_{z}^{*}$ is similar to $D^{*}$ using the result of Shields, Theorem 2.7. We now consider an example of a Hilbert space with a non-diagonal reproducing kernel. The proof, being similar to that of Example 4.11, is omitted.
@@
Let $M_{z}^{*}$ be the backward shift operator on a Hilbert space $\mathcal{H}_{K}$ with reproducing kernel
@@
We close this section with a proposition stating that to a Hermitian holomorphic complex bundle with a tensor product structure satisfying some given conditions, there corresponds a Cowen-Douglas operator. This result establishes the operator realization of the tensor product of holomorphic complex bundles.
@@
Let $\mathcal{E}$ be a Hermitian holomorphic line bundle over $\mathbb{D}$ and let $\gamma$ be a non-vanishing holomorphic cross-section of $\mathcal{E}$. Suppose that $E=\bigvee\limits_{w\in\mathbb{D}}\gamma(w)$ and $\{e_{i}\}_{i=0}^{\infty}$ is an orthonormal basis for $E$. If
@@
for $\gamma(w)=\sum\limits_{i=0}^{\infty}\gamma_{i}(w)e_{i}$ and some constant $\delta>0$, then there exists an operator $T\in\mathcal{B}_{1}(\mathbb{D})$ such that $\mathcal{E}_{T}=\mathcal{E}_{D_{\alpha}^{*}}\otimes\mathcal{E}$ and $T\sim_{s}D_{\alpha}^{*}.$
@@
Recall that the Cowen-Douglas conjecture states that for operators $T$ and $S$ in $\mathcal{B}_{1}(\mathbb{D})$ having $\overline{\mathbb{D}}$ as a $K$-spectral set, $T$ is similar to $S$ if and only if $\lim\limits_{|w|\rightarrow 1}\mathcal{K}_{T}(w)/\mathcal{K}_{S}(w)=1$. In [10, 11], Clark and Misra constructed examples to show that the Cowen-Douglas conjecture is false and claimed that the ratio of the metrics can be a better indication of similarity than the ratio of the curvatures. In particular, several examples are provided in Theorem 1 of [11] to support this claim. Theorem 3.4 and Proposition 3.6 in this note offer further evidence.
Every Cowen-Douglas operator is unitarily equivalent to the adjoint of a multiplication operator on a reproducing kernel Hilbert space consisting of holomorphic functions. For the Hardy and weighted Bergman spaces, Theorem 1 in [11], Theorem 3.8 in [55], Theorem 3.4 in [25], and Theorem 3.4 and Proposition 3.6 of this paper suggest that, under certain conditions, the ratio of the metrics is a similarity invariant of the backward shift operator on these spaces. In particular, for an $n$-hypercontraction $T\in\mathcal{B}_{1}(\mathbb{D})$, from Theorem 0.1 in [38], Theorem 2.4 in [16], Theorem 2.3 in [24], and Theorem 3.4 of this paper, we conclude that $T$ is similar to $S_{n}^{*}$ on the reproducing kernel Hilbert space $\mathcal{M}_{n}$ if and only if $a_{w}:=h_{\mathcal{E}_{T}}(w)/h_{\mathcal{E}_{S_{n}^{*}}}(w)$ is bounded and bounded away from 00 if and only if the curvature $\mathcal{K}_{\mathcal{E}_{D_{T}}}$ of the holomorphic line bundle $\mathcal{E}_{D_{T}}$ induced by the defect operator $D_{T}$ satisfies $\mathcal{K}_{\mathcal{E}_{D_{T}}}=-\frac{\partial^{2}}{\partial w\partial\overline{w}}\varphi(w)$ for some bounded subharmonic function $\varphi$ defined on $\mathbb{D}$.
Clark and Misra constructed contractive backward shift operators in 2.5 of [11], such that the ratio of the metrics of the holomorphic line bundles associated with the operators is bounded and bounded away from 00 that are not similar to each other. This indicates that the ratio of the metrics is not a similarity invariant for all backward shift operators on Hilbert spaces with diagonal reproducing kernels. One can consider a follow-up question.